### v2

# Cycling Race Time Prediction: A Personalized Machine Learning Approach Using Route Topology and Training Load

Abstract
Predicting cycling duration for a given route is essential for training planning and event preparation. Existing solutions rely on physics-based models that require extensive parameterization, including aerodynamic drag coefficients and real-time wind forecasts, parameters impractical for most amateur cyclists. This work presents a machine learning approach that predicts ride duration using route topology features combined with the athlete’s current fitness state derived from training load metrics. The model learns athlete-specific performance patterns from historical data, substituting complex physical measurements with historical performance proxies. We evaluate the approach using a single-athlete dataset (N=96 rides) in an N-of-1 study design. After rigorous feature engineering to eliminate data leakage, we find that Lasso regression with Topology + Fitness features achieves MAE=6.60 minutes and R²=0.922. Notably, integrating fitness metrics (Chronic Training Load [CTL], Acute Training Load [ATL]) reduces error by 14% compared to topology alone (MAE=7.66 min), demonstrating that physiological state meaningfully constrains performance even in self-paced efforts. Progressive checkpoint predictions enable dynamic race planning as route difficulty becomes apparent.
1.  Introduction
$1.$  Motivation
Cycling duration prediction is a practical problem faced by athletes at all levels. Whether planning a training session, estimating arrival times for a group ride, or pacing a competitive event, knowing how long a route will take provides valuable information for decision-making. This need becomes particularly relevant when routes involve significant elevation changes, where naive distance-based estimates fail to account for the time cost of climbing.
Current solutions to this problem fall into two categories. Physics-based tools, such as Best Bike Split, solve equations of motion considering power output, aerodynamic drag, rolling resistance, and environmental conditions. While accurate, these approaches require parameters that are difficult to obtain: aerodynamic drag coefficients typically require wind tunnel testing or field measurements, and accurate wind forecasts are rarely available for arbitrary routes. The alternative, simple rule-of-thumb estimates based on average speed, ignores the significant impact of terrain and the athlete’s current form.
A gap exists for an accessible prediction method that accounts for route difficulty and individual fitness without requiring specialized measurements. Machine learning offers a potential solution: by learning from an athlete’s historical performance data, a model can implicitly capture the relationship between route characteristics, fitness state, and resulting duration.
1.2.  Objectives
The main objective of this work is to develop a machine learning model capable of predicting cycling duration from route topology features and athlete fitness state.
The specific objectives are:
  1. 1.
    Combine route topology features with training load metrics as model inputs.
  2. 2.
    Demonstrate the feasibility of personalized prediction using single-athlete historical data in an N-of-1 study design.
  3. 3.
    Evaluate checkpoint-based progressive predictions that update estimates as the ride progresses.
  4. 4.
    Identify the most predictive features through importance analysis.
1.3.  N-of-1 Study Design Rationale
This work employs an N-of-1 study design, analyzing 96 rides from a single athlete. This approach is justified by three factors: (1) High inter-individual variability in cycling physiology—power profiles, VO2max, lactate thresholds, and anaerobic capacity vary 2-5x across amateur cyclists. Population models would require extensive physiological testing (laboratory VO2max, critical power profiling) impractical for the target use case. (2) Personalized prediction goal—we seek to answer “How long will this route take ME?” rather than “How long for an average cyclist?” Personalized models learn athlete-specific pacing patterns and strengths. (3) Data availability—modern cyclists generate thousands of GPS activities, providing abundant within-subject data without requiring sparse multi-athlete cohorts. This design has proven effective in chronobiology, nutrition, and clinical medicine for personalized prediction tasks with high inter-subject variability.
1.4.  Paper Structure
The remainder of this paper is organized as follows. Section 2 reviews related work in cycling performance prediction, machine learning for sports analytics, and training load modeling. Section 3 describes the methodology, including data collection, feature engineering, and model architecture. Section 4 presents the experimental evaluation and results. Section 5 demonstrates practical application through a case study. Section 6 discusses the findings, limitations, and directions for future work.
2.  Related Work
2.1.  Cycling Performance Prediction
The prediction of cycling performance has traditionally relied on physics-based models. Di Prampero et al. [6] established the foundational equation of motion for cyclists, relating power output to resisting forces: aerodynamic drag, rolling resistance, gravitational force on gradients, and drivetrain losses. Martin et al. [12] validated this mathematical model for road cycling power, demonstrating accurate prediction of power requirements under controlled conditions. This equation forms the basis of commercial tools that predict race times and optimal pacing strategies.
Physics-based approaches achieve high accuracy when properly parameterized. However, they require inputs that are difficult to obtain in practice. The aerodynamic drag coefficient (CdA) depends on rider position, equipment, and clothing; measuring it accurately requires wind tunnel testing or specialized field protocols. Rolling resistance varies with tire pressure, surface conditions, and temperature. Wind speed and direction along a route are rarely available with sufficient spatial and temporal resolution.
Critical Power (CP) models offer an alternative framework, characterizing an athlete’s capacity through two parameters: CP (the power sustainable indefinitely) and W’ (the finite work capacity above CP). These models predict time to exhaustion at given intensities but do not directly address the route-specific duration prediction problem, as they assume constant power output.
$2.$  Machine Learning in Sports Analytics
Machine learning has been applied extensively to sports prediction problems [5]. In cycling, researchers have used gradient boosting methods to predict race outcomes, leveraging historical results, course profiles, and athlete rankings. Deep learning approaches, including LSTM networks, have modeled sequential performance data to capture temporal patterns in athlete form. Jobson et al. [7] demonstrated the utility of cycling training data analysis for performance modeling, while Menaspà et al. [13] characterized the physical demands of professional cycling using power meter data.
A notable distinction exists between predicting rankings and predicting actual times. Most published work focuses on classification or ranking tasks: determining which athlete will finish first or estimating finishing position. Direct time prediction receives less attention, likely due to the additional complexity of modeling absolute performance rather than relative performance.
For running, the TRAP framework demonstrated checkpoint-based progressive prediction, where estimates are refined as an athlete passes intermediate timing points. This approach has not been systematically applied to cycling, where checkpoint infrastructure is less common and routes are more variable.
2.3.  Training Load and Fitness Modeling
The relationship between training and performance has been modeled through fitness-fatigue frameworks. The impulse-response model represents an athlete’s performance potential as the difference between a positive fitness component and a negative fatigue component, both responding to training stimuli with different time constants.
Training Stress Score quantifies the physiological load of individual workouts based on intensity and duration relative to threshold power [3]. Chronic Training Load (CTL) represents the exponentially weighted average of Training Stress Score over approximately 42 days, serving as a proxy for fitness. Acute Training Load (ATL) uses a shorter time constant (approximately 7 days) to capture recent fatigue. Training Stress Balance (TSB), calculated as CTL minus ATL, indicates an athlete’s freshness and theoretical readiness to perform.
These metrics are widely used in training prescription and have been incorporated into performance prediction models. However, their integration with route-specific features for duration prediction remains underexplored.
2.4.  Research Gap and Contribution
Existing approaches to cycling duration prediction either require extensive parameterization (physics-based models) or focus on ranking rather than absolute time prediction (machine learning approaches). The combination of route topology features with training load metrics for direct duration prediction represents an unexplored area.
This work addresses the gap by proposing an accessible machine learning approach that:
  • Uses readily available data: route files (GPX) and training history from standard cycling computers and platforms.
  • Learns athlete-specific performance patterns without requiring aerodynamic or physiological testing.
  • Incorporates fitness state through established training load metrics.
  • Supports checkpoint-based progressive prediction for real-time updates during rides.
The N-of-1 study design acknowledges that individual variation in cycling performance makes personalized models more practical than population-level approaches for time prediction.
3.  Methodology
3.1.  Data Sources
The data for this study come from a single amateur cyclist’s training records, collected over multiple years through standard cycling equipment. All data are sourced from Intervals.icu, a training analysis platform that aggregates data from cycling computers and power meters.
3.1.1. Data Platform
Raw data are processed through a medallion data architecture. The ingestion layer (bronze) retrieves data from the Intervals.icu API; a cleaning pipeline (silver) applies type coercion, schema validation, and unit standardization. The platform runs daily, providing clean, consistently formatted data for analysis.
Data cleaning filters activities to ensure quality and relevance: (1) outdoor rides only (indoor trainer sessions with zero GPS movement excluded), (2) minimum duration of 30 minutes to exclude short commutes and equipment tests, (3) complete GPS and altitude streams (activities with >>10% missing data points excluded), and (4) rides with sufficient effort (excluding recovery spins below 0.5 IF). These criteria ensure the model learns from representative training and racing efforts rather than casual rides. The final dataset contains 96 activities spanning 31 months (March 2023–October 2025).
3.1.2. Activities
The activities dataset contains metadata for each recorded ride: start time, elapsed time, moving time (time spent in motion, excluding stops), distance, and activity type. Rides are distinguished from other activity types (runs, swims, strength training) through the type field. Each activity has a unique identifier that links to the corresponding sensor streams. Moving time serves as the prediction target throughout this work.
3.1.3. Wellness and Training Load
Daily wellness records capture physiological state: body weight, CTL, ATL, and TSB. These are computed using established formulas based on historical power data and reflect the athlete’s state at the start of each day.
3.1.4. Activity Streams
Each activity includes second-by-second sensor data: GPS coordinates, altitude, speed, power, heart rate, and cadence. Route topology features are derived from altitude and GPS streams; performance metrics from power and speed data.
3.2.  Feature Engineering
Table 1 summarizes all 79 engineered features with their formulas.
Category Feature Formula Description
Basic Metrics (7)
total_distance Route distance (km)
total_ascent $\sum\max(0,\Delta h)$ Cumulative elevation gain (m)
total_descent $\sum\max(0,-\Delta h)$ Cumulative elevation loss (m)
elevation_min/max/avg Elevation statistics (m)
elevation_gain_per_km $G/d$ Ascent normalized by distance (m/km)
Gradient Variability (3)
punchiness_score $\sigma(|\Delta g|)$ Std of gradient changes
gradient_std $\sigma(g)$ Standard deviation of gradient
gradient_cv $\sigma(g)/|\bar{g}|$ Coefficient of variation
Climb Detection (9)
num_climbs count Detected climbs (ClimbPro algorithm)
num_hc/cat1/…/cat4 Climbs by category (HC $\geq$80k, …, Cat4 $\geq$8k)
total_climb_score $\sum d_{i}\times g_{i}$ Sum of Garmin scores (Eq. 1)
max_climb_score $\max(d_{i}\times g_{i})$ Hardest individual climb
climb_density $n_{\text{climbs}}/d$ Climbs per km
Climb Characteristics (4)
avg_climb_gradient $\sum(g_{i}d_{i})/\sum d_{i}$ Distance-weighted avg gradient (%)
avg/max/total_climb_length Climb distance metrics (m)
Gradient Distribution (5)
pct_slope_* $d_{\text{bucket}}/d$ % route in 6 gradient buckets
pct_above_5/8/10% $d_{>x}/d$ % route above threshold
Technical Features (6)
[…]
turn_density $n_{\text{turns}}/d$ Sharp turns per km
recovery_distance $\sum d_{i}\cdot\mathbf{1}_{|g|<2\%}$ Flat distance after climbs (m)
technical_descent $\sum d_{i}\cdot\mathbf{1}_{g<-5\%,\Delta\theta>45^{\circ}}$ Steep descent + turn combos (m)
max_sustained_gradient $\max(\bar{g}_{500\text{m}})$ Max 500m rolling avg gradient (%)
longest_climb_distance Longest detected climb (m)
Athlete Fitness (4)
ctl $\text{EMA}_{42}(\text{TSS})$ Chronic Training Load (fitness)
atl $\text{EMA}_{7}(\text{TSS})$ Acute Training Load (fatigue)
tsb $\text{CTL}-\text{ATL}$ Training Stress Balance (form)
ramp_rate $\Delta\text{CTL}/\Delta t$ Fitness rate of change
Rolling Zone Hours (48)
rolling_*d_power_z*_hours $H_{w,z}$ (Eq. 6) 7 power zones $\times$ 4 windows
rolling_*d_hr_z*_hours $H_{w,z}$ (Eq. 6) 5 heart rate zones $\times$ 4 windows
Total 79 features
Table 1: Complete feature set: route topology (27) and athlete state (52). Notation: $d$ = distance, $g$ = gradient, $h$ = altitude, $\theta$ = bearing, $\Delta$ = difference, $\sigma$ = std dev, $\mathbf{1}$ = indicator.
Figure 1: Elevation profile of La Desértica MTB race (110 km, 2443m ascent). Top: altitude with 13 detected climbs (ClimbPro: $\geq$3%, $\geq$500m) highlighted in red. Bottom: gradient distribution showing sections above 5% (orange) and 8% (red).
The model inputs are organized into two categories: route topology features that characterize the physical demands of a course, and athlete state features that capture training history. Table 1 summarizes the feature groups.
3.2.1. Route Topology Features
Route topology features quantify difficulty from GPS and altitude data. Beyond basic metrics (distance, elevation), we derive features capturing terrain nuances. Figure 1 illustrates these features on an example route.
Climb Detection
Following Garmin ClimbPro’s algorithm, climbs are continuous segments where gradient exceeds 3% for at least 500m. Each is scored as:
$S_{\mathrm{climb}}=d\times g$ (1)
where $d$ is segment length in meters and $g$ is average gradient in percentage points (e.g., 8 for 8%), yielding an HC score of 80,000 for a 10km climb at 8%. Scores above 80,000 are Hors Catégorie (HC); 64k, 32k, 16k, 8k for Cat 1–4; >>1,500 for uncategorized climbs. We additionally compute Tour de France scores ($g^{2}\times d$) as a separate feature. From detected climbs we derive: count by category (HC, Cat 1–4), total and maximum scores, average gradient weighted by distance, climb density (climbs per km), and length metrics (average, maximum, total climb distance). Short climbs (1–5 min) tax anaerobic capacity; sustained climbs (>>10 min) depend on aerobic power [10].
Gradient Variability
Figure 2: Punchiness score analysis. Top: gradient profile with overall route punchiness $P$ (standard deviation of gradient changes). Bottom: rolling punchiness over 1km windows for visualization, identifying sections with highest gradient irregularity. The model uses the global $P$ as a single feature; the rolling view illustrates where variability concentrates.
Variable terrain causes 26% greater fatigue than constant grades at equal average power [9]. This occurs because gradient variability forces repeated anaerobic efforts during accelerations on steep sections, prevents steady-state metabolic adaptation, and requires constant power adjustments that increase neuromuscular fatigue. We compute a punchiness score:
$P=\text{std}(\Delta g)\quad\text{where}\quad\Delta g_{t}=|g_{t+1}-g_{t}|$ (2)
The punchiness score $P$ is the standard deviation of absolute gradient changes across the entire route, yielding a single scalar that captures terrain irregularity. Unlike gradient standard deviation (which measures overall spread), punchiness captures the frequency and magnitude of changes, making it more sensitive to rolling terrain that appears flat by traditional metrics. Higher values indicate rolling terrain with repeated accelerations. Figure 2 visualizes local 1km rolling windows to show where variability is highest, while the model uses the global standard deviation as a feature.
Figure 3: Punchiness score comparison across three routes with similar distance and elevation but different gradient variability. High punchiness (left) indicates frequent gradient changes requiring repeated power surges; low punchiness (right) shows steady grades allowing constant pacing. This metric captures difficulty beyond total elevation.
Recovery and Technicality
After significant climbs, flat sections allow W (anaerobic work capacity) reconstitution with time constants of 377–580s [17]. We compute recovery distance as flat terrain within 500m of climb ends:
$D_{\text{recovery}}=\sum_{i\in\mathcal{R}}d_{i}\cdot\mathbf{1}_{|g_{i}|<2\%}$ (3)
where $\mathcal{R}$ is the set of points within 500m after climbs with $g_{\text{avg}}>3\%$. Routes with minimal recovery distance accumulate fatigue faster as W cannot reconstitute between efforts.
Technical descents are identified where gradient $5\%$ coincides with bearing change $45^{\circ}$:
$D_{\text{tech}}=\sum d_{i}\cdot\mathbf{1}_{g_{i}<-5\%}\cdot\mathbf{1}_{\Delta\theta_{i}>45^{\circ}}$ (4)
These sections require sustained attention despite low power output, limiting recovery [1].
Figure 4: Recovery distance analysis. Climbs (red, $g>3\%$) and recovery zones (green, $|g|<2\%$ within 500m of climb ends). Flat sections after climbs allow W reconstitution; routes with minimal recovery distance accumulate fatigue faster.
Gradient Distribution
Physiological response varies non-linearly with gradient, with steep sections (¿6%) requiring anaerobic contributions and moderate gradients (2-6%) sustainable aerobically. We compute distance percentages in gradient buckets (negative, 0-2%, 2-4%, 4-6%, 6-10%, 10%+) and thresholds (% route above 5%, 8%, 10%). Figure 5 illustrates gradient analysis on an example route.
Figure 5: Gradient distribution analysis showing three complementary views: histogram of gradient values, cumulative distance in gradient buckets, and sustained gradient over rolling 500m windows. Routes with similar total elevation can have vastly different gradient profiles, affecting physiological demand.
Figure 6: Climb categorization for dataset routes using Garmin ClimbPro algorithm. Most routes contain Cat 3-4 climbs (8-32k score), while challenging routes feature HC and Cat 1 climbs (>>64k score). Climb category distribution captures route difficulty beyond simple elevation gain.
Figure 7: Maximum sustained gradient analysis. The 500m section with highest average gradient (peak physiological demand) is highlighted on the elevation profile, showing location of hardest sustained effort within the route.
Elevation Profile Shape
We also compute route distance in six gradient buckets (descending, flat, false flat, moderate, hard, very hard) and percentage of total ascent in each route third. Back-loaded routes may induce greater fatigue from glycogen depletion.
Sharp Turns and Turn Density
Sharp turns require braking, gear changes, and power surges to accelerate out, disrupting rhythm and accumulating fatigue [14]. We detect sharp turns where bearing change exceeds $45^{\circ}$:
$\Delta\theta_{t}=\min\bigl(|\theta_{t+1}-\theta_{t}|,\,360^{\circ}-|\theta_{t+1}-\theta_{t}|\bigr)$ (5)
where $\theta_{t}$ is the smoothed bearing (5-point rolling mean to reduce GPS jitter). A turn is sharp when $\Delta\theta>45^{\circ}$. Turn density normalizes by distance: $T_{d}=n_{\text{turns}}/d_{\text{km}}$.
Figure 8: Sharp turn detection examples. Each panel shows the route geometry around a detected turn (red dot at origin) where bearing change exceeds $45^{\circ}$. Axes show distance in meters from turn point; green/orange markers indicate travel direction.
Additional Route Metrics
Elevation gain per km normalizes climbing by distance, enabling comparison across routes of different lengths.
3.2.2. Athlete State Features
Training intensity distribution predicts performance [15]. Following Banister’s impulse-response model [4] and Seiler’s time-in-zone methodology [16], we compute rolling zone hours:
$H_{w,z}=\frac{1}{3600}\sum_{t-w<t^{\prime}<t}s_{z,t^{\prime}}$ (6)
where $H_{w,z}$ is hours in zone $z$ over window $w$ days, and $s_{z,t^{\prime}}$ is seconds in zone $z$ on day $t^{\prime}$. Strict inequality prevents data leakage.
Windows: 7d (acute), 14d (short-term), 30d (monthly), 60d (chronic)
Power zones: Z0–Z6 (7 zones) $\times$ 4 windows = 28 features
Heart rate zones: Z0–Z4 (5 zones) $\times$ 4 windows = 20 features
High hours in upper zones over short windows indicate recent fatigue-inducing intensity; accumulated endurance zone hours over longer windows reflect aerobic base.
3.2.3. Training vs. Deployment Feature Sets
Critical distinction: This work evaluates features for training but deploys models using only prediction-time available features.
During model development, we evaluated three feature configurations to determine what information drives duration prediction:
  1. 1.
    Topology-only (27 features): Route characteristics from GPS and altitude streams—distance, elevation, climbs, gradient distribution, punchiness. All extractable from GPX files before a ride.
  2. 2.
    Topology + Fitness (31 features): Adds athlete state at ride start—CTL, ATL, TSB, ramp rate. Requires historical training data.
  3. 3.
    Topology + Fitness + Zones (79 features): Adds rolling zone hours (7d, 14d, 30d, 60d windows across power and heart rate zones). Captures recent training load distribution.
Results: The best model (Lasso Topology + Fitness) achieves MAE=6.60 min and R²=0.922 using 31 features (27 topology + 4 fitness). Fitness metrics improve accuracy by 14% over topology alone (MAE=7.66 min), demonstrating that physiological state meaningfully constrains performance even in self-paced efforts.
Deployment scenarios:
  • Best accuracy: Use Topology + Fitness model (31 features) when historical training data are available. Requires athlete’s current CTL, ATL, TSB, and ramp rate from prior weeks of training.
  • Cold start / Universal: Use Topology-only model (27 features) when predicting for new athletes without training history, or when analyzing routes without fitness context. Accepts any GPX file (GPS + elevation) with slightly reduced accuracy (MAE=7.66 vs 6.60 min).
  • What-If scenario analysis: The Topology + Fitness model enables projections like “Given this route and my target fitness in 12 months (CTL=65, TSB=+10), what duration should I expect?” This is possible because fitness features use historical training load, not ride performance results.
$3.$  Progressive Prediction Framework
For pre-ride planning, the model predicts duration from the complete route topology. For race-day applications, we evaluate checkpoint-based progressive predictions where estimates are updated as the athlete completes portions of the route.
Checkpoint Definition
Checkpoints are defined at fixed distance percentages (25%, 50%, 75%, 100%) along the route. At each checkpoint, topology features are recomputed from the GPS track up to that point, providing a cumulative route profile up to the current position.
Feature Recalculation
All topology features (climbs, gradients, punchiness) are recalculated on the truncated track. For example, at 50% completion on a 100km route, features are extracted from the first 50km only. This creates a prediction based on terrain encountered so far, revealing route difficulty as it unfolds.
Use Case
Progressive predictions answer: “Given the terrain I’ve ridden so far, how much total time will this route take?” This differs from real-time remaining time prediction, instead providing updated full-route duration estimates as more terrain is revealed. Prediction stability depends on route difficulty distribution—front-loaded climbs yield stable early estimates; back-loaded routes show greater prediction evolution.
3.4.  Model Selection and Validation
3.4.1. Model Candidates
We evaluate three model families: (1) Regularized linear models (Ridge, Lasso, ElasticNet) with L1/L2 penalties to prevent overfitting, (2) Random Forest with Bayesian hyperparameter optimization via Optuna [2], and (3) Baseline models (mean, median, simple linear regression on distance + elevation) to establish prediction difficulty.
3.4.2. Nested Cross-Validation
With $N=96$ rides and up to 79 features (topology + fitness + zones), overfitting risk is severe. We employ nested cross-validation to obtain unbiased performance estimates:
Outer Loop (Evaluation)
5-fold stratified CV by distance quintiles ensures test folds remain isolated, never used for hyperparameter selection. Stratification by distance prevents imbalanced folds (long rides concentrated in one fold).
Inner Loop (Hyperparameter Tuning)
For each outer training fold:
  • Linear models: 3-fold CV with automatic alpha selection (RidgeCV, LassoCV)
  • Random Forest: 3-fold CV + Optuna Bayesian optimization (30 trials per fold)
Leakage Prevention
  1. 1.
    Feature imputation (median strategy) fitted on training fold only
  2. 2.
    Standardization (z-score) fitted on training fold only
  3. 3.
    Rolling zone features use strict temporal inequality (Equation 6)
3.4.3. Evaluation Metrics
Models are compared by: (1) MAE (Mean Absolute Error in minutes), interpretable for athletes, (2) , proportion of variance explained, (3) Train-Test Gap, difference between train and test MAE detecting overfitting (gap <2<-2 min indicates severe overfitting), and (4) CV Stability, coefficient of variation across folds indicating robustness.
4.  Model Evaluation
4.1.  Dataset
The dataset comprises 96 outdoor cycling activities recorded between March 2023 and October 2025 from a single amateur cyclist (male, Functional Threshold Power 232W). Activities were classified as free rides (N=92) or races (N=4) using duration deviation from expected time. Table 2 summarizes route and performance characteristics.
Metric Mean $\pm$ SD Range
Distance (km) 24.4 $\pm$ 13.1 [10.2, 110.1]
Elevation (m) 520 $\pm$ 339 [100, 2,575]
Moving Time (min) 114.1 $\pm$ 57.5 [55.9, 546.9]
Avg Gradient (%) 2.1 $\pm$ 0.8 [0.5, 4.7]
Climbs per Ride 3.2 $\pm$ 2.8 [0, 13]
Table 2: Dataset statistics for 96 cycling activities.
4.1.1. Fitness State Distribution
Training load metrics exhibit moderate temporal variation (Figure 9): CTL ranges from 0 (training restart after break) to 63.5 (peak fitness), with CV=54.3%. TSB ranges from 58-58 (deeply fatigued) to +15.8+15.8 (well-rested), with CV=89.5%. The presence of zero values in early 2023 indicates return from training break, while sustained CTL >30>30 in late 2023-2024 reflects consistent training volume.
Figure 9: Fitness metric temporal analysis. Top: CTL/ATL/TSB time series showing moderate variation (CV $\approx$ 54%) over 31-month period. Middle: Distribution histograms. Bottom: Model residuals vs. CTL showing weak correlation (r=-0.37), suggesting fitness metrics do not strongly explain prediction errors.
4.1.2. Dataset Diversity
The dataset spans diverse route types with different characteristics: short punchy climbs, long steady grades, and mixed terrain. Routes with similar distances can have vastly different duration and difficulty due to gradient distribution, climb categories, and punchiness scores.
4.1.3. Feature Engineering
From the 96 activities, we extracted 79 features: 27 route topology metrics (Table 1), 4 fitness metrics, and 48 rolling zone hours. Figure 10 visualizes the relationship between distance, elevation, and duration.
Figure 10: Dataset scatter plot showing relationship between distance, elevation gain, and moving time. Point size indicates number of climbs. Strong correlation visible between distance/elevation and duration (R²=0.908 for simple linear model), but significant variance remains unexplained without terrain complexity features.
Figure 11: Feature correlation matrix for top topology features. Distance and total ascent show strong correlation (r=0.82), while punchiness and climb density capture orthogonal route characteristics. Low correlation between many features justifies inclusion of diverse topology metrics.
4.2.  Model Comparison
Nested cross-validation results reveal that incorporating fitness state is essential for accurate prediction. The Lasso (Topology + Fitness) model achieves the best performance with MAE=6.60 minutes and R²=0.922. This represents a significant improvement over the Topology-only baseline (MAE=7.66 min). Complex non-linear models (Random Forest) perform poorly (MAE=13.36 min), suffering from severe overfitting due to the small sample size (N=96), validating the choice of regularized linear models for this domain.
Figure 12: Model performance comparison showing MAE and R² across all models. Lasso (Topology + Fitness) achieves lowest error. Random Forest shows high variance and severe overfitting despite Optuna hyperparameter tuning. Error bars show standard deviation across 5 CV folds.
Model MAE (min) Status
Baseline (Mean) 31.82 $\pm$ 8.04 0.43-0.43 $\pm$ 0.71
Linear (Dist+Elev) 6.91 $\pm$ 1.02 0.908 $\pm$ 0.077 Baseline
Lasso (Topo+Fit) 6.60 $\pm$ 1.13 0.922 $\pm$ 0.061 Best
ElasticNet (Topo+Fit) 6.68 $\pm$ 1.01 0.909 $\pm$ 0.093 Good
Lasso (Topo+Fit+Zones) 6.72 $\pm$ 1.05 0.917 $\pm$ 0.069 Good
Ridge (Topo+Fit) 6.80 $\pm$ 0.85 0.907 $\pm$ 0.094 Good
ElasticNet (Topo) 7.29 $\pm$ 0.51 0.889 $\pm$ 0.104 OK
Ridge (Topo) 7.46 $\pm$ 0.42 0.888 $\pm$ 0.101 OK
Lasso (Topo) 7.66 $\pm$ 1.84 0.898 $\pm$ 0.070 Baseline
Random Forest (All) 13.36 $\pm$ 6.52 0.679 $\pm$ 0.211 Overfit
Table 3: Model performance via 5-fold nested cross-validation. MAE and R² reported as mean $\pm$ standard deviation across folds. Adding fitness features improves accuracy by approximately 14% over topology alone.
4.2.1. Fitness Metrics Improve Prediction Accuracy
[…]
4.2.2. Regularization and Feature Engineering
The disparity between Lasso (MAE=6.60) and Random Forest (MAE=13.36) highlights the importance of regularization in N-of-1 studies where N is small relative to P features. By penalizing non-predictive coefficients, Lasso effectively selects a sparse subset of impactful features, whereas Random Forest memorizes noise in the training folds. Additionally, the removal of result-dependent features (such as VAM or average speed) ensures the model is deployment-ready. The reported accuracy relies solely on data available prior to the ride: the GPX file and the athlete’s training history. Figure 13 shows that validation error plateaus around 60 samples, confirming dataset size adequacy for linear models.
Figure 13: Learning curve for Lasso (Topology) model. Validation error (orange) plateaus at $\sim$60 samples with small train-validation gap, confirming effective regularization and adequate dataset size. Shaded regions show standard deviation across CV folds.
4.2.3. Prediction Accuracy
Figure 14 shows predicted vs. actual durations for the best model (Lasso Topology + Fitness). Points cluster tightly around the diagonal (perfect prediction line) with R²=0.922. Residuals show no systematic bias across the duration range.
Figure 14: Predicted vs. actual moving time for Lasso (Topology + Fitness) model. Points cluster near diagonal (perfect prediction), with mean absolute error of 6.60 minutes and R²=0.922 across 96 activities. No systematic bias visible across duration range (55-547 minutes).
4.3.  Error Analysis
Figure 15 breaks down prediction errors by route difficulty. The model performs consistently across short/long and flat/hilly routes, with no systematic bias. Larger errors (>>10 min) occur on extreme routes (very long or very steep), but remain proportionally small relative to total duration.
Figure 15: Error analysis by route characteristics. Model performs consistently across difficulty levels with no systematic bias. Larger absolute errors on extreme routes remain proportionally small (MAPE << 5% even for longest/steepest routes).
Figure 16: Model performance stratified by route difficulty tiers (distance and elevation). MAE remains stable across easy, moderate, hard, and very hard routes, demonstrating robustness to varying route characteristics.
$4.$  Feature Importance
To understand which route characteristics drive predictions, we analyze feature importance using SHAP (SHapley Additive exPlanations) [11] values on the best model (Lasso with topology features).
4.4.1. Global Importance
Figure 17 shows mean absolute SHAP values, quantifying each feature’s average contribution to predictions. The top 5 features account for over 80% of predictive power:
  1. 1.
    Distance (22.5 min): Primary driver of duration
  2. 2.
    Total Ascent (16.5 min): Cumulative vertical meters add time non-linearly
  3. 3.
    Elevation Gain/km (12.5 min): Normalized difficulty metric (m/km)
  4. 4.
    % Grade 4-6% (5.1 min): Percent of route in moderate climbing zone
  5. 5.
    Punchiness Score (4.1 min): Terrain variability affects pacing
Figure 17: SHAP feature importance for Lasso (Topology) model. Bars show mean —SHAP value— in minutes, indicating average contribution to prediction. Distance and elevation metrics dominate, but terrain complexity features (punchiness, climb density) contribute 25-30% of predictive power.
4.4.2. Feature Value Effects
Figure 18 visualizes how feature values influence predictions: Distance and ascent show strong linear positive relationships (higher values $\rightarrow$ longer time), while punchiness score exhibits non-linear effects—high punchiness (rolling terrain) adds 5-10 minutes even at equal distance/elevation, validating the importance of gradient variability in duration prediction.
Figure 18: SHAP summary plot showing feature value effects. Each point represents one activity, colored by feature value (red=high, blue=low). Horizontal position shows SHAP value (impact on prediction). Distance and ascent show strong positive effects; punchiness adds time for high values.
5.  Application: Track 101 MTB Case Study
Having validated the model’s accuracy, we now demonstrate its practical application on a real-world route planning scenario. We evaluate the model on Track 101 MTB 2025, a challenging $\sim$101 km mountainous route with 3,279m elevation gain (32.5 m/km difficulty index, 1.5x dataset mean).
5.1.  Progressive Checkpoint Predictions
Predictions are made at 25%, 50%, 75%, and 100% distance checkpoints using Lasso (Topology) trained on the full 96-activity dataset.
Prediction Methodology
At each checkpoint, the model predicts total route duration using topology features extracted from the GPX up to that point. At 25% completion, only the first quarter of the route is analyzed; at 100%, the complete profile is used. This reveals how duration estimates evolve as terrain difficulty becomes apparent.
The model uses only the 27 topology features extracted from the truncated GPX file (start to checkpoint). No fitness state, no historical data. This simulates pre-ride planning where an athlete progressively examines route segments.
For Track 101 MTB 2025, we created checkpoint GPX files at 25%, 50%, 75%, and 100% distance and predicted total duration at each stage. The substantial increase from 201 min (25%) to 648 min (100%) reveals this route’s back-loaded difficulty—the first quarter appears manageable, but cumulative climbing in later sections dramatically extends the predicted finish time.
5.1.1. Prediction Evolution
Table 4 shows predictions evolve substantially as terrain is revealed. The prediction more than triples from 25% completion (201 min) to 100% (648 min), indicating this route’s difficulty is heavily back-loaded.
Checkpoint Dist (km) Elev (m) Climbs Predicted Time (min)
25% 25.2 671 5 200.9
50% 50.4 1,023 8 293.6
75% 75.6 1,939 13 450.7
100% 100.7 3,279 18 647.6
Table 4: Progressive checkpoint predictions for Track 101 MTB 2025. Prediction more than triples from 25% to 100% due to back-loaded climbing (only 5 climbs in first quarter, 18 total).
Figure 19: Progressive checkpoint predictions for Track 101 MTB 2025. Top-left: Predicted duration more than triples from 25% to 100% completion. Top-right: Distance and elevation accumulation through route. Bottom-left: Prediction change rate accelerates in final quarter (7.87 min/% vs. 3.71 min/% in first quarter). Bottom-right: Only 5 climbs detected at 25%, revealing back-loaded difficulty distribution.
5.1.2. Route Profile and Checkpoint Visualization
Figure 20 shows the elevation profile of Track 101 MTB 2025 with checkpoint locations and predicted durations marked. The profile reveals the back-loaded difficulty: relatively gradual terrain in the first half gives way to sustained climbing in the final 50km, with major ascents concentrated between km 60–90.
Figure 20: Track 101 MTB 2025 elevation profile with progressive checkpoint predictions. Vertical lines mark 25%, 50%, 75%, and 100% distance checkpoints. Each annotation shows cumulative distance and predicted duration (e.g., 3h20m at 25%, 10h47m at finish). The profile reveals back-loaded climbing: gentle rolling terrain until km 50, followed by sustained climbs in the final half. Total route: 100.7 km, 3,279m ascent, 18 climbs.
5.1.3. Back-Loaded Route Difficulty
The route accumulates 18 climbs total, but only 5 are visible at 25% completion (Figure 19). Prediction change rate accelerates from 3.71 min/% (25-50%) to 7.87 min/% (75-100%), indicating steeper, longer climbs concentrate in the final quarter. With a difficulty index of 32.5 m/km (vs. dataset mean of 21.0 m/km), this route represents a “very challenging” classification—critical information for nutrition and pacing strategy.
5.2.  ClimbPro Analysis
Applying the Garmin ClimbPro algorithm to Track 101 MTB reveals how climb difficulty accumulates across checkpoints. Figure 21 shows the elevation profile with all detected climbs highlighted by category, and the gradient distribution below. Table 5 shows the progressive detection of climbs as the route unfolds.
Figure 21: ClimbPro analysis for Track 101 MTB 2025. Top: Elevation profile with 18 detected climbs highlighted by category (Cat 3 in orange, Cat 4 in yellow, uncategorized in green). Bottom: Gradient distribution showing sections above 8% (red), 5–8% (orange), and below 5% (green). The route contains 6 Cat 3 climbs concentrated in the final 40km.
Checkpoint Climbs Cat 3 Cat 4 Score Max 500m Grad Longest (km)
25% 5 1 0 47,244 12.02% 3.83
50% 8 1 0 61,117 12.02% 3.83
75% 13 3 1 130,740 18.93% 6.01
100% 18 6 1 215,879 19.13% 6.01
Table 5: ClimbPro analysis for Track 101 MTB 2025. Climb count and category distribution reveal back-loaded difficulty: 72% of total climb score (156k of 216k) concentrates in the final 50% of the route. Cat 3 climbs increase from 1 to 6 between 50% and 100%.
The ClimbPro scoring ($S=d\times g$) quantifies this progression: at 50% distance, only 28% of total climb difficulty has been encountered (61k of 216k score). The sudden jump at 75% (131k score) signals the route’s demanding final quarter.
5.3.  Maximum Sustained Gradient Analysis
The max sustained gradient over 500m metric identifies the single hardest sustained effort on the route—the 500-meter section with highest average gradient. Figure 22 highlights this section on Track 101 MTB.
Figure 22: Maximum sustained gradient analysis for Track 101 MTB 2025. Top: Full route with the hardest 500m section highlighted in red (19.6% average gradient at km 73.8–74.3). Bottom: Zoomed view of the max effort section showing 99m elevation gain over 500m. This metric captures peak physiological demand—the single hardest sustained effort regardless of total elevation.
For Track 101 MTB, the maximum sustained gradient is 19.6% located at km 73.8–74.3, within the demanding final quarter. The evolution across checkpoints (12.0% at 25%, 12.0% at 50%, 18.9% at 75%, 19.6% at 100%) confirms the back-loaded difficulty.
This metric captures peak physiological demand beyond what aggregate statistics reveal. A route with identical total elevation could have maximum sustained gradients of 8% (steady) or 20% (punchy)—requiring fundamentally different pacing strategies. The 19.6% maximum on Track 101 indicates sections requiring near-maximal effort, even for well-trained athletes, and validates the “very challenging” classification.
Practical Application
Progressive predictions enable race-day decisions: at 50% completion (50.4 km), the model predicts 293.6 min total, revealing that the second half will take significantly longer than the first. Athletes can adjust effort accordingly, conserving energy for the demanding final quarter. This “checkpoint intelligence” transforms static pre-race duration estimates into dynamic race-day tools.
5.4.  Fitness Feature Contribution
The Lasso (Topology + Fitness) model achieves 14% lower MAE than topology alone, confirming that fitness state contributes predictive value. However, interpreting individual fitness coefficients requires caution: with N=96 and correlated predictors, coefficient directions may reflect training data patterns (e.g., fitter athletes attempting harder routes) rather than isolated causal effects.
For Track 101, varying fitness parameters while holding topology constant yields predictions within a 5-minute range, confirming that route topology dominates duration prediction. The minimum predicted duration remains $\sim$10h50m regardless of fitness assumptions, revealing that the 32.5 m/km difficulty index and 18 detected climbs fundamentally constrain achievable speed.
The practical implication is clear: for challenging routes like Track 101, athletes should prioritize route-specific preparation—pacing strategy, nutrition planning, and checkpoint targeting—rather than expecting large time gains from fitness optimization alone.
6.  Discussion and Conclusion
6.1.  Discussion
6.1.1. Topology Sets the Baseline, Fitness Refines the Prediction
Route topology is the primary driver of duration, explaining over 88% of the variance in our dataset. Features like punchiness (gradient variability) capture difficulty beyond simple elevation gain [9]. However, the integration of fitness metrics (CTL, ATL) provides a critical 14% improvement in accuracy (MAE reduced from 7.66 to 6.60 min).
This finding suggests that while athletes may self-pace to manage effort, their “cruising speed” is statistically associated with their chronic training load (CTL) and acute fatigue state (ATL). The 14% error reduction (MAE improved by $\approx$1 minute) may seem modest, but compounds meaningfully over training cycles: more accurate predictions enable better pacing strategy and event preparation.
6.1.2. Fitness Improves Prediction, but Interpretation Requires Caution
While fitness features improve predictive accuracy by 14%, interpreting individual coefficients for “what-if” scenario analysis requires caution. With N=96 and correlated predictors, learned coefficients may reflect confounding patterns in the training data—fitter athletes tend to attempt harder routes—rather than isolated causal effects. For example, the model’s fitness coefficients cannot reliably answer questions like “how much faster will I be if I taper before race day?”
Despite this limitation, the aggregate contribution of fitness features is valid: including CTL, ATL, and related metrics consistently reduce prediction error across cross-validation folds. Future work with larger datasets or experimental designs (e.g., repeated measurements under controlled fitness variations) could enable more reliable causal inference for training periodization guidance.
6.1.3. The Importance of Rigorous Feature Selection
Our evaluation highlights the risk of data leakage in sports analytics. Initial iterations using VAM (Vertical Ascent Speed) yielded artificially high accuracy (R²>>0.97) but failed to generalize for future predictions. By strictly isolating a priori features (Topology and Fitness) from a posteriori results (Speed, VAM), we achieved a realistic and deployable model (R²=0.922).
6.1.4. Regularization Critical for Small Datasets
With N=96 and 27-79 features, overfitting risk is severe. Random Forest, despite aggressive Bayesian hyperparameter tuning via Optuna (30 trials per fold), achieves only MAE=13.36—worse than simple baselines—with catastrophic overfitting. Lasso’s L1 penalty proves essential, producing stable predictions and automatic feature selection. This negative result has practical value: practitioners should default to regularized linear models for small sports performance datasets until N >> 500.
6.1.5. N-of-1 Study Design Justification
Single-athlete studies are appropriate when: (1) High inter-individual variability: Cycling performance depends on Functional Threshold Power, VO2max, lactate threshold, W’, body composition, and skill—parameters varying 2-5x across amateur cyclists. Population models would require extensive physiological testing (lab VO2max, power profiling) impractical for the target use case. (2) Personalized prediction goal: This work targets individual athletes seeking “How long will this route take ME?” rather than “How long for an average cyclist?” Personalized models learn athlete-specific strengths (climbing vs. flat speed) and pacing tendencies. (3) Data availability: Modern cyclists generate thousands of GPS activities, providing abundant within-subject data without requiring sparse multi-athlete cohorts.
Similar N-of-1 designs have proven effective in chronobiology, nutrition, and clinical medicine [8]. The key is sufficient within-subject variation in predictor variables (route characteristics) and outcome (duration), which our dataset provides.
6.2.  Limitations
Dataset Size
While N=96 is adequate for Lasso regression (learning curve plateaus at $\sim$60 samples), it precludes exploration of deep learning models. Gradient boosting or neural networks might capture non-linearities inaccessible to linear models, but require N >> 500 for stable training on this feature dimensionality.
Single Athlete
Results may not generalize to cyclists with different physiology, skill levels, or pacing strategies. A competitive racer might show strong fitness-performance coupling (no self-pacing), while a novice might exhibit different duration-gradient relationships. The model is intentionally personalized, requiring individual training data.
Environmental Factors Excluded
Weather conditions (wind, temperature, precipitation), road surface quality, and traffic are not captured. Wind alone can alter duration by 10-30% on exposed routes. GPS data lack environmental sensors; integrating weather API data are future work.
Pacing Assumptions
The model assumes consistent pacing strategy (self-selected for training, maximal for races). Strategic pacing variations (conserving energy for final climb) are not modeled. Duration prediction implicitly averages over typical pacing behavior learned from historical data.
GPS Accuracy
Elevation data from barometric altimeters have $\pm$5m precision, while GPS elevation can drift $\pm$20m. Systematic errors in climb detection or gradient calculation could propagate to predictions, though high-quality devices mitigate this limitation.
6.3.  Future Work
Multi-Athlete Validation
Expanding to N=10-50 athletes would enable population-level models and quantify personalization benefit magnitude. Transfer learning could pre-train on population data, then fine-tune to individuals with limited personal history, reducing cold-start requirements for new users.
Environmental Integration
Weather API integration (wind speed/direction, temperature) would improve race-day prediction accuracy. Wind affects duration by 10-30% on exposed routes; historical weather data aligned to GPS tracks could train weather-aware models for improved real-world applicability.
Real-Time Pacing Guidance
Current work predicts total duration; real-time systems could provide live feedback: “You’re 2 min ahead of predicted pace, ease off to avoid bonking on final climb.” This requires modeling remaining time given current position and accumulated fatigue, an open research problem in endurance sports analytics.
Segment-Level Predictions
Rather than total duration, predicting time for individual climbs or segments enables Strava segment performance forecasting and climb-specific race preparation. This requires segment-level feature extraction and residual fatigue modeling across cumulative efforts.
6.4.  Conclusion
This work demonstrates that machine learning models can predict cycling duration from route topology and fitness state with high accuracy (MAE=6.60 minutes, R²=0.922) using only consumer-grade GPS data and historical training load. Unlike physics-based approaches requiring aerodynamic testing and real-time wind forecasts, our method learns athlete-specific performance patterns from historical rides, making personalized predictions accessible to amateur cyclists.
Key contributions include: (1) Novel terrain features grounded in exercise physiology (punchiness, climb detection matching Garmin ClimbPro, gradient distribution), (2) Rigorous feature selection eliminating data leakage from result-based metrics (VAM, speed, power), ensuring deployment validity, (3) Regularization importance demonstrated through Random Forest overfitting on N=96, validating Lasso’s robustness for small sports datasets, and (4) Fitness integration showing training load metrics (CTL, ATL) provide 14% accuracy improvement over topology alone, though individual coefficient interpretation requires caution due to predictor correlation.
The N-of-1 study design proves viable for personalized sports analytics, leveraging abundant within-subject GPS data to overcome inter-individual variability challenges. Progressive checkpoint predictions enable race planning, though back-loaded route difficulty limits early estimate stability.
This approach enables practical applications without specialized equipment: pre-ride duration estimates for training planning, race pacing strategy, and route difficulty assessment. The work showcases an end-to-end machine learning pipeline suitable for demonstrating data collection, rigorous feature engineering, proper validation, and deployment-ready prediction capabilities.
6.5.  Data and Code Availability
The feature extraction pipeline, model training code, and analysis notebooks are available at https://github.com/fran-aguila/health-hub. Due to privacy considerations, the raw GPS activity data and training logs are not publicly released. Researchers interested in replicating this work can apply the methodology to their own cycling data using standard export formats from platforms such as Strava, Garmin Connect, or Intervals.icu.